The rms speed of the molecules in 1.0 g of hydrogen gas is1800 m/s.
(a) What is the total translational kinetic energy of the gasmolecules?
(b) What the thermal energy of the gas?
(c) 500 J of work are done to compress the gas while, in thesame process, 1200 J of heat energy are transferred from the gas tothe environment. Afterward, what is the rms speed of themolecules?
Answer #1
rms speed v = 1800 m/s, mass m = 1.0 g, mole mass M = 2.0g/mol, mole number n = m/M = 0.5
a) the total translational kinetic energyof the gas molecules
Et = mv2/2 = 1.0*10-3 *18002/2 = 1620 J = 3nRT/(2M)
b) the thermal energy of thegas
E = 5nRT/(2M) = (5/3) *[3nRT/(2M)] = (5/3)*Et = 2700J
c) Q = 1200 J, W = 500 J, final temperature = T', thermal energy =E', rms speed = v'
ΔE = W - Q = -700 J = E' - E
so E' = E + ΔE = 2000 J
E' = 5nRT'/(2M) = (5/3)*Et' = (5/3)mv'2/2
v' = √(6E'/5m) = 1549 m/s